Ticket Number: HEA-NOC/20090813-7 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20090813 10:48 UTC+1 Problem Start: 20090812 17:50 UTC+1
Ticket Closed: 20090814 18:05 UTC+1 Problem End: 20090812 17:51 UTC+1
Site/line: St. Angela's
There appears to have been a very brief (approximately 20 seconds) flap of
the BT circuit HEAN107521 to STACS.
Brief loss of connectivity to/from STACS.
20090813 10:45 LT Circuit has been stable since the flap but have asked BT
to carry out an alarm check and we are awaiting the results. BT ref:
20090813 13:05 BMA BT rang, confirmed that there were transmission alarms
on their SDH equipment/network, but couldn't give an exact cause because
they cleared before investigation.
20090814 18:05 LT Circuit has been stable ever since. Closing ticket.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20090813-7
HEAnet Limited. Registered in Ireland, No. 275301.